### v1

# Enhancing Large Language Models for Text-to-Testcase Generation

Abstract
Context: Test-driven development (TDD) is a widely employed software development practice that involves developing test cases based on requirements prior to writing the code. Although various methods for automated test case generation have been proposed, they are not specifically tailored for TDD, where requirements instead of code serve as input.
Objective: In this paper, we introduce a text-to-testcase generation approach based on a large language model (GPT-3.5) that is fine-tuned on our curated dataset with an effective prompt design.
Method: Our approach involves enhancing the capabilities of basic GPT-3.5 for text-to-testcase generation task that is fine-tuned on our curated dataset with an effective prompting design. We evaluated the effectiveness of our approach using a span of five large-scale open-source software projects.
Results: Our approach generated 7k test cases for open source projects, achieving 78.5% syntactic correctness, 67.09% requirement alignment, and 61.7% code coverage, which substantially outperforms all other LLMs (basic GPT-3.5, Bloom, and CodeT5). In addition, our ablation study demonstrates the substantial performance improvement of the fine-tuning and prompting components of the GPT-3.5 model.
Conclusions: These findings lead us to conclude that fine-tuning and prompting should be considered in the future when building a language model for the text-to-testcase generation task.
keywords:
TDD, Text-to-Testcase generation.
1 Introduction
Test-Driven Development (TDD) is an agile software development methodology that prioritizes creating test cases before the actual coding process (Beck, 2022). This methodology is structured around iterative cycles that involve the development of test cases based on a given requirement, the implementation of code to pass these test cases, and then refactoring to continuously improve the code. TDD is claimed to significantly improve code quality and handle requirements better than ’more traditional, heavy weight predictive methods’ (Wasmus and Gross, 2007). The implementation of TDD led to a reduction in defects by approximately 50%, in comparison to a similar system developed with ad-hoc unit testing methods (Maximilien and Williams, 2003). Despite these benefits, one of the primary drawbacks of this methodology is the time and effort required to write test cases. The initial investment in creating comprehensive and effective test cases is substantial and represents the largest overhead cost in the TDD process (Beck, 2022). However, this investment is a critical component of TDD.
[…]
The Large Language Models (LLMs), being advanced language models, are capable of performing various Natural Language Processing (NLP) tasks such as text generation, question answering, and classification. Specifically in the field of software engineering (SE), LLMs are increasingly being applied for tasks such as requirements engineering, code generation, vulnerability identification, test generation, and software maintenance. The use of LLMs for generating test cases from natural language descriptions of software methods (text-to-test case) is a relatively new application.
In this paper, we introduce a Text-to-Test generation approach, which involves a fine tuning component for the basic GPT-3.5 using our dataset specifically curated from open-source Java projects and an effective prompting design. The model fine-tuning is aimed to teach the LLM the mapping between the method description (text), and the corresponding test (testcase). Once the basic GPT-3.5 model is fine-tuned, during the model inference phase, an unseen method description (i.e., the method description is referred to as the requirement) with the designed prompt is used as input into the fine-tuned GPT-3.5 to generate the corresponding test case without requiring the code. Crucially, our approach differs from traditional methods by generating test cases directly from method descriptions without requiring the implemented code (method). This allows for the generation of test cases in the absence of actual code, a significant shift from existing practices.
We compare our approach with basic GPT-3.5 (OpenAI, 2023), Bloom (Luccioni et al., 2022), and CodeT5 (Wang et al., 2021) fine-tuned on our newly curated benchmark dataset, consisting of 163k pairs of method descriptions and test cases. We developed an evaluation dataset with 7K test methods that span five large-scale open source software projects Apache Commons Lang (Lang, 2022), JFreeChart (JFreeChart, 2022), Apache Common CLI (CLI, 2022), Apache Common CSV (CSV, 2022), Google Gson (Gson, 2019). For each model, we evaluate the quality of the generated test cases in three aspects: syntax correctness, requirement alignments, and code coverage. In summary, our objective is to address the following research questions (RQs):
RQ1. How effective is the proposed approach for the Text-to-Test generation task?
Results
.  The proposed approach generates test cases that are 78.5% syntactically correct and 67.09% aligned with the requirements while achieving a code coverage of 61.7%, which substantially outperforms the baselines (basic GPT-3.5-turbo, Bloom, and CodeT5).
RQ2. What is the effect of the fine-tuning step for the LLM-based text-to-testcase generation task?
Results
.  The proposed fine-tuning component improves syntax correctness by 223%, requirements alignment by 164%, and code coverage by 153%.
RQ3. What is the effect of the prompting step for the LLM-based text-to-testcase generation task?
Results
.  The proposed prompt design improves the syntax correctness by 124%, the requirements alignment by 82%, and the code coverage by 58%.
RQ4. What types of errors occur in the incorrectly generated test cases?
Results
.  We observed that 78.5% of the generated test cases are syntactically correct, while 21.5% of the generated test cases are incorrect. Out of the incorrectly generated test cases (21.5%) we observed that 11.3% were associated with assertion errors, 2.4% with value errors, 6.9% with syntax errors, and the remaining 0.9% with other types of errors.
Based on our results, we draw the following recommendations:
  • 1.
    Recommendation 1: The basic usage of ChatGPT may not produce a desirable performance for the text-to-testcase generation task. Our RQ1 shows that the basic usage of ChatGPT (GPT-3.5-turbo with no fine-tuning and a basic prompt) may not achieve a desirable performance as we observe that it performs worse than other LLMs like Bloom.
  • 2.
    Recommendation 2: Both fine-tuning and effective prompting design should always be considered when using a large language model. Our RQ1 shows that our proposed approach performs the best for the text-to-testcase generation task, while our RQ2 and RQ3 confirm the substantial benefits of the fine-tuning and the prompting component. However, both components should be considered together, as the results show that both fine-tuning and effective prompting designs outperform LLMs with one individual component alone.
Novelty & Contributions.
  1. 1.
    We present a new Text-to-Test generation approach based on a fine-tuned large language model with an effective prompting design.
  2. 2.
    The ablation study confirms the substantial benefits of fine-tuning and prompt design for the text-to-test generation task.
Open Science. To support open science, we publish the studied dataset, scripts (i.e., data processing and model fine-tuning), and experimental results in a GitHub repository. (https://github.com/LLMforTDD/LLMforTDD).
Structure. Section 2 discusses background and related work. Section 3 presents our approach. Section 4 presents our experiment. Section 5 presents the results. Section 6 discloses the threats to validity, and Section 7 concludes the paper.
2 Background and Related Work
In this section, we provide background knowledge of Test-Driven Development (TDD) and Large Language Models (LLMs) and discuss related work to motivate the paper.
2.1 Test Driven Development (TDD)
Test Driven Development (TDD) (Beck, 2022; Astels, 2003) is a software development practice that suggests writing test cases before implementing source code. Starting with a comprehensive requirement, TDD follows an iterative process of providing a small unit test case that aligns with the requirement while developing a small chunk of source code to pass the test case at a time. The practice aligns with the Agile Development Methodology that emphasizes fast continuous feedback loops and enhances collaboration between team members. Recent research found that TDD can increase the understanding of requirements and produce a higher number of test cases than the traditional code-first method (Mueller and Zeller, 2002; Janzen, 2008; Erdogmus and Arisan, 2005). Despite the numerous benefits discussed in the literature, TDD is still manual, time-consuming, and labor-intensive, highlighting the need for automated Text-To-Testcase generation.
2.2 Large Language Models (LLMs)
A language model (LM) is a statistical model trained to predict the next word in a sequence (Jurafsky and Martin, 2020), e.g., for the sequence “Paris is the capital of”, it will predict with a very high likelihood that “France” is the next word in the sequence. Large language models (LLMs) are LMs with a substantially large number of weights and parameters and a complex training architecture trained to perform various downstream NLP tasks, e.g., text generation, question answering, and text classification (Wei et al., 2022). The recent advancements in LLMs, most notably with chatbot platforms built on LLMs, e.g., OpenAI’s ChatGPT, have led to several applications in software engineering (SE) (Nguyen-Duc et al., 2023; Hou et al., 2023). These include, among others, requirements engineering automation tasks (Arora et al., 2023), code generation and completion (Fan et al., 2023), software vulnerability (Fu et al., 2023), test generation (Yuan et al., 2023), and software maintenance (Nguyen-Duc et al., 2023).
The use of LLMs for generating test cases from text (text-to-test case) is a relatively new application. It involves LLMs understanding a natural language (NL) description of the software method and then generating appropriate test cases that can be used to verify that the software meets its specifications. While there is interest in this area, it still remains underexplored due to the complexity of accurately interpreting descriptions and generating meaningful and comprehensive test cases. However, as LLMs become more advanced, their ability to perform such specialized tasks will likely improve, making text-to-testcase generation a promising area for future research and application in software engineering.
Prompting. In generative AI models (built on LLMs), the input is provided as a natural language instruction to elicit a desired response or output (Brown et al., 2020). The input is known as a “prompt”. How well a prompt is crafted determines the quality and relevance of the AI’s output. There are different prompting strategies, and one typically needs to experiment with various prompts to elicit accurate responses or outputs from the models (Nguyen-Duc et al., 2023).
Fine-tuning In the context of Generative AI and ML refers to taking a pre-trained model and further training it on a specific, usually smaller, dataset to adapt it to a particular task or set of tasks. This process is prevalent in applications where a model, initially trained on a diverse and general dataset, needs to be specialized for more specific purposes. This specialized dataset is tailored to meet the particular needs or context of the intended application or task. This approach is beneficial as it leverages the extensive learning from the initial training phase, making the fine-tuning process more efficient than training a new model from scratch on the specialized data. Therefore, fine-tuning enhances the model’s performance and accuracy in tasks requiring a more focused understanding or specialized knowledge (Brown et al., 2020).
2.3 Related Work
In this section, we discuss related work on automated test case generation, LLM in software engineering, and LLM for test case generation to highlight the significance of our paper.
2.3.1 Automated Test Case Generation
Automated test case generation is an emerging technique that leverages neural networks and other deep learning architectures to automatically create test cases for software applications. This approach aims to enhance the efficiency and effectiveness of software testing by automating the generation of test inputs and expected outputs. Researchers leveraged various approaches for automated test case generation, e.g., randomization, heuristics, search-based algorithms, and deep learning.
Random-based test case generation approaches like Randoop Pacheco and Ernst (2007) are proposed to automatically generate test cases randomly. Randoop uses a feedback-directed random test generation algorithm. It starts by generating a sequence of method calls and this sequence is generated randomly. Randoop uses feedback from previous tests to guide the selection process. For example, Randoop may be more likely to generate a sequence of method calls that includes methods that have not been tested in the past. It observes the results of the execution to determine whether the program behaved as expected and creates test cases based on the results of the execution.
Search-based test case generation approaches like Evosuite Fraser and Arcuri (2011) uses an evolutionary algorithm to generate test cases that maximize code coverage. It works by first generating a random population of test cases. Then, it executes the test cases and measures the code coverage. The test cases that achieve the highest code coverage are selected to be the parents of the next generation of test cases. EvoSuite mutates the parents to create the next generation of test cases. The mutations are designed to increase the code coverage of the test cases. It repeats this process until it reaches a time limit or until it finds a set of test cases that achieve a certain level of code coverage.
[…]
2.3.2 LLMs for Test Case Generation
Large Language Models (LLMs) have emerged as powerful tools in various fields of natural language understanding and generation, including software engineering (Hou et al., 2023). Notably, Hou et al. (Hou et al., 2023) and Wang et al. (Wang et al., 2023) found that LLMs have been widely used to support various software testing tasks. For example, Schafer et al. (Schäfer et al., 2023) proposed TestPilot, a LLM-based test generation tool for JavaScript that automatically generates unit tests for all API functions in an npm package. Siddiq et al. (Siddiq et al., 2023) leveraged Large Language Models (LLMs) for test case generation. Similarly, Yuan et al. (Yuan et al., 2023) and Xie et al. (Xie et al., 2023) leveraged ChatGPT for test case generation.
[…]
3 Enhancing Large Language Models for Text-To-Test Generation
Problem Formulation. In this paper, we propose a text-to-testcase generation approach based on a fine-tuned LLM with an effective prompting design. In particular, we formulate the text-to-testcase as a generation task using a fine-tuned LLM (GPT-3.5-turbo). The model is fine-tuned on a large dataset, and is mined and curated from open-source Java projects. During the model fine-tuning, our LLM aims to learn the mapping between the method description (text), which is the comments or documentation that describe the purpose, functionality, and usage of a method, and the corresponding test case (testcase) which is a test method that includes a set of assert statements designed to verify whether the implemented method (method) is aligned with the given requirement or not (i.e., functionally correct). Once fine-tuned, during the model inference phase, an unseen method description is used as input into the fine-tuned LLM to generate the corresponding test cases without requiring the implemented method (method), which differs from most existing test case generation approaches. Our approach consists of three phases: (1) model fine-tuning, (2) prompt design, and (3) model inference. Below, we describe the details of our approach (Figure 1).
Figure 1: An overview of our approach (i.e., the prompting-based fine-tuning GPT-3.5-turbo model for the Text-To-Testcase generation task.
3.1 Model Fine-tuning
While LLMs demonstrate impressive capabilities in understanding and generating natural language (Hou et al., 2023), they may lack the nuanced contextual understanding required for specialized applications, e.g., the text-to-testcase generation task. The fine-tuning of LLMs is essential to enhance their performance and adaptability to our text-to-testcase generation task. Thus, it is important to have a representative and well-labeled dataset for the specific domain to achieve effective fine-tuning results. However, the existing test case generation datasets are not suitable for our task. For example, Tufano et al. (Tufano et al., 2022) introduced the Method2Test dataset, which is a pair of <<<Method,Testcase>>>, but lacking the corresponding method description that is required for our approach. Thus, we need to carefully curate the fine-tuning dataset.
3.1.1 Dataset Curation
For the fine-tuning of the model, we started from the Methods2Test dataset (Tufano et al., 2022). This dataset consists of 780K pairs of Java focal methods and their corresponding test cases, collected from 91K open-source Java projects. Since the method description is unavailable, we extended the Methods2Test dataset. Starting from the 91K open-source Java projects, we cloned each project from the provided GitHub URL and performed the following steps:
(Step 1) Collecting Code Structural Information. The structural information of the code, such as method and class names are very important as we used to identify the relationship between the focal methods (i.e., methods to be tested) and their corresponding test methods. To collect such information, we used the tree-sitter parser (Tree-sitter, 2022) to analyze the structure of source code within the given project based on the detailed abstract syntax tree (AST) representation of source code. With this step, we are able to identify the list of method and class names within the project.
(Step 2) Identify Test Classes. With the parser, we are able to identify the test classes (i.e., the classes specifically created to test the functionality of other classes in a program). Thus, any class that contains at least one method with the @Test annotation is identified as a test class. Such the @Test annotation in Java is normally used to signal that the annotated method should be executed as a test case.
(Step 3) Identify Focal Classes. For each test class in Step 2, we aim to identify the focal class associated with the given test class. To do so, we employ the following two heuristics:
File Path Matching: Following the best practices of JUnit testing, they suggest placing focal classes and corresponding test classes in a mirrored directory structure. For example, if a class is located at src/main/java/Foo.java, the corresponding JUnit test cases should be placed in the test class file src/test/java/FooTest.java.
File Name Matching: The name of a test class is usually composed of the name of the focal class, starting with a ”Test” prefix or suffix. For example, the test case for the class Foo.java would probably be named as FooTest.java. Thus, we can identify the test classes and their corresponding focal classes with the heuristics of file path and file name matching.
(Step 4) Identify Test Methods and Focal Methods. We identify the test methods for each test class based on the @Test annotation. Similar to Step 3, we apply the method name matching heuristic (i.e., removing Test prefix or suffix in the method name) to identify the corresponding focal method and its method description.
With these steps, we construct the dataset <<<Text,Testcase,Method>>> in the format of triplets. The dataset consists of 163k pairs of JUnit test cases and corresponding method description. We note that the methods will not be used for model fine-tuning like prior work, but are only used for evaluating test cases.
3.1.2 Fine-Tuning
When performing model fine-tuning, each training example generally consists of a single input (i.e., method description) example and its associated output (i.e., test case) without giving detailed instructions or including multiple examples in the same prompt normally used for the base models (GPT-3.5). Following the OpenAI’s legacy fine-tuning best practices [^1] https://platform.openai.com/docs/guides/legacy-fine-tuning, each training example typically consists of a single input example and its associated output, which is typically constructed in the following format: {"input": "<Text>", "output": "<test case>"}. We extract method descriptions and corresponding test cases into a list of prompts and completions for fine-tuning. We initiate a fine-tuning job using OpenAI’s CreateFineTuneJob.create method, specifying the training data and model to fine-tune (GPT-3.5-turbo). We use ChatCompletion.create to generate test cases with the fine-tuned model, providing: The model engine ID from the fine-tuning job, the input is the method description prompt. The output is completion of the generated test case.The template includes the {"prompt": "<Text>", "completion": "<test case>". The input is a single input-output pair. We used the curated dataset constructed in Section 3.1.1 for model fine-tuning (the pairs of <text, test case>). Once completed, we use the fine-tuned model for generating test cases in the next step.
Figure 2: Structure of our Improved prompt and Basic prompt.
Figure 3: Model input and the generated test case and the ground-truth for evaluation.
3.2 Model Inference
Once the tuning is complete, we use the fine-tuned model to perform inference tasks. To apply our approach using the fine-tuned model, we require input in the form of a prompt, with the textual description. We note that, in our study for inference, we developed an evaluation dataset (Section 4.2) with 7K test methods that span across five large-scale open-source software projects (i.e., Lang, Chart, Cli, Csv, Gson). We used the java parser to identify the public methods and we extracted the method description of each public method for all the 5 projects. The average cost associated with fine-tuning GPT-3.5 is \$0.0057 per test case and \$40.28 in total for all 5 evaluation projects.
3.2.1 Test Case Generation
We require the fine-tuned model and the prompt as input for test case generation. The model then generates a response containing the generated test cases.
For test case generation, the most important step was to develop a prompt that works well, as the prompts provide context and guidance to the model. Without a (clear) prompt, the model would not know what information or response is expected. The process of formulating prompts for software engineering tasks is inherently subjective, and there are no established empirical guidelines for optimizing prompt effectiveness and robustness. We conducted several experiments involving various prompt combinations, ranging from basic to customized prompts, to address concerns associated with prompt optimization and devised our improved prompt. We present two prompts below. A basic prompt (our starting point) and an improved prompt (after several iterations).
Basic prompt It is a straightforward instruction for LLM that does not include additional instructions, context, or specialized formatting. A generic or minimal input relies on the model’s capabilities to generate a response based on the provided context. We use the basic prompt as a context in the model. The structure of the basic prompt is as follows:
"/* Write a junit test case /**Description **/"
Improved prompt It is the iterated version based on multi-step optimizations. We tested the prompt in the first step on a few samples from a small validation set (not from our curated or evaluation set to avoid data leakage), manually analyzed the generation performance, and optimized the generation results by providing ChatGPT with a series of new prompts.
As in Figure 2, we designed the prompt to instruct the LLM with a more specific context. It emphasizes the importance of proper and relevant assert statements, avoiding repetition, following a structured test case format, and adding the context details for a more structured test case. A detailed breakdown of the prompt template components and their intended purpose is:
  • 1.
    Introductory Comment: Instructs the model to focus on creating a test case, highlighting the need for relevant assert statements and avoiding repetition.
  • 2.
    Test Case Template: Provides a skeleton for the test method structure, including the @Test annotation and placeholders for common assert statements.
  • 3.
    Class and Method Name: Specifies the target class and method name for which the test is to be written, ensuring the generated test is relevant.
  • 4.
    Method Description: Provides a description of the method, giving context to generate a test case.
Once the test cases are generated, we conduct a series of checks and corrections in post-processing to ensure the validity of these generated test cases.
3.2.2 Post-Processing
When extracting tests from LLM’s response, syntactic errors may occur because the tokens limit is exceeded, resulting in truncated tests. We thus perform a three-part post-processing on the generated test cases.
  1. 1.
    Verifying Test Signatures: We inspect the beginning of each generated test case to ensure the presence of four specific keywords in sequence: @Test, public, void, and test[MethodName]. This sequence is crucial for the proper syntax of JUnit test methods.
  2. 2.
    Adding Missing Keywords: If any of these keywords are missing in the initial sequence of the generated test method, we insert them accordingly. This step is essential to guarantee that the test methods can be executed successfully.
  3. 3.
    Correcting Incomplete Parentheses: We noticed that some of the generated test cases contain incomplete parentheses, such as () or . To address this issue, we perform corrections on these long sequence test cases, ensuring that all parentheses are correctly balanced.
These measures ensure that the generated test methods align with the input method descriptions and adhere to the syntactical requirements for successful execution.
4 Experimental Design
In this section, we present the research questions and the experimental design.
4.1 Research Questions
Our text-to-testcase generation approach involves two key components (fine-tuning and prompting). Thus, our experiment aims to investigate the performance of our text-to-testcase generation approach when compared to other baseline models, conduct an ablation study to investigate the benefits of fine-tuning and prompting steps, and investigate the types of errors that occur in the incorrectly generated test cases. Therefore, we formulate the following research questions.
  1. RQ1
    How effective is the proposed approach for the Text-to-Test generation task?
  2. RQ2
    What is the effect of the fine-tuning step for the LLM-based text-to-testcase generation task?
  3. RQ3
    What is the effect of the prompting step for the LLM-based text-to-testcase generation task?
  4. RQ4
    What types of errors occur in the incorrectly generated test cases?
Table 1: A Summary of the Evaluation Dataset for the LLM-based text-to-testcase generation models.
Project Abbr. Domain Version # LOC # Description
Commons-Lang Lang Utility 3.12.0 56,787 1878
Commons-Cli Cli Cmd-line interface 1.5.0 36,289 368
Commons-Csv Csv Data processing 1.10.0 25,485 31
Gson Gson Serialization 2.10.1 24,686 174
Jfreechart Chart Visualization 1.5.4 152,686 4599
4.2 Evaluation Dataset
To mitigate the risk of data leakage between model fine-tuning and model evaluation, we decided to curate a new benchmark dataset for evaluating the LLMs for the text-to-testcase generation task. Table 1 provides a summary of the studied five open-source Java projects namely, Apache Commons Lang (Lang, 2022), JFreeChart (JFreeChart, 2022), Apache Common CLI (CLI, 2022), Apache Common CSV (CSV, 2022), Google Gson (Gson, 2019). We selected these five projects due to the following reasons. We chose open-source software projects to support the reproducibility and replicability of our work, allowing researchers and practitioners to reproduce experiments and validate our results. We chose diverse domains and varying sizes of projects to ensure that the results are not bound to specific contexts.
4.3 Baseline Comparisons
To fairly evaluate our proposed approach, we chose the other three competitive LLMs (i.e., GPT-3.5, Bloom and CodeT5) as a baseline comparison. We choose the following three LLMs based on the various model size, types of model architectures, open/closed source, and fit in our GPU infrastructure. Table 2 shows the analysis of the characteristics of the baseline models.
Model Model Size Architecture Open/Closed Source Fit the GPU?
GPT-3.5 175B Decoder only Closed Source No
Bloom 560M Decoder only Open Source Yes
CodeT5 770M Encoder-Decoder Open Source Yes
Table 2: Comparison of Different Models
To investigate the benefits of the fine-tuning and prompting steps for our approach, we conduct an ablation study by selectively removing or ablating one of the components and observing the impact on the overall performance of our approach. Thus, we conducted a comprehensive experiment of the four variations (i.e., With and Without Finetune, and Basic and Advanced Prompt) on three LLMs (i.e., GPT-3.5-turbo, Bloom, CodeT5), totaling 12 models. Since we observed that the fine-tuning and the prompting for Bloom and Code T5 perform best, for the remainder of the paper, we will focus on the comparison of our approach (GPT-3.5-turbo FineTune + Imprv.P), i.e., the prompting-based fine-tuning GPT-3.5-turbo model with the following baseline.
  • 1.
    GPT-3.5-turbo (FineTune + Basic.P)
  • 2.
    GPT-3.5-turbo (NoFineTune + Imprv.P)
  • 3.
    GPT-3.5-turbo (NoFineTune + Basic.P)
  • 4.
    Bloom (Fine-Tune + Imprv.P)
  • 5.
    CodeT5 (Fine-Tune + Imprv.P)
4.4 Hyper-Parameter Settings for Fine-Tuning.
We fine-tune the two models, namely CodeT5 (Wang et al., 2021) and Bloom (Luccioni et al., 2022), using the following settings. We run our experiment on two NVIDIA GeForce RTX 3090 GPUs with 24 GB vRAM, an Intel(R) Core(TM) i9-9980XE CPU@ 3.00GHz with 36 core processors, and 64G RAM. The hyperparameter settings for fine-tuning are reported in Table 3.
Hyperparameter Value
Learning Rate 2e-5
Warmup Steps 1000
Weight Decay 0.01
Batch Size 2
Gradient Accumulation Steps 4
Learning Rate Scheduler Type inverse_sqrt
Epoch 20
Table 3: Hyper-parameters and their values for CodeT5 and Bloom
Figure 4: (RQ1)An example of test cases generated by GPT-3.5-turbo, and other baseline approaches
4.5 Evaluation Measures
The LLMs take ”Improved prompt and text” as input and the focal method as ground truth for evaluating the generated test cases (Figure 3). We evaluate the quality of the test cases generated by LLMs based on the following three aspects:
  • 1.
    Syntax Correctness measures the percentage of the generated test cases that are syntactically correct without any compiler errors, which is a prerequisite for the TDD process. The given test case is considered syntactically correct if it can compile successfully.
  • 2.
    Requirements Alignment measures the percentage of generated test cases that match the specified requirement. Since the ground truth code represents the correct implementation of the method requirement, any test cases that execute successfully against the ground truth code are considered correct (i.e., they correctly align with the given requirement and invoke the appropriate functions to be tested with the correct test inputs and expected outputs). Therefore, we define that a test case is aligned with a given requirement if it can be executed with the ground truth code. To assess this, we execute the combination of the ground-truth code and the generated test cases.
  • 3.
    Code Coverage measures the completeness of the generated test cases for a given description. Ideally, in addition to aligning with the requirements, the test cases must be completed (i.e. executing all lines of the ground-truth code solution). Thus, we measure the line-level code coverage of the generated test cases using the Jacoco (Library, 2019).
5 Experimental Results
In this section, we present the results with respect to our four research questions.
RQ1: How effective is the proposed approach for the Text-to-Test generation task?
Figure 5: The experimental results of the ablation study with six different models
Results. Our proposed approach (i.e., the fine-tuned GPT-3.5-turbo with an effective prompting design) outperforms all other LLMs (basic GPT-3.5-turbo, Bloom, and CodeT5). Figure 5 presents the results of our approach and the baseline models for a text-to-testcase generation task. We find that our proposed approach (i.e., the fine-tuned GPT-3.5-turbo with an effective prompting design) is able to generate test cases that are 78.5% being syntactically correct, 67.09% being aligned with the requirements while achieving a code coverage of 61.7%, which substantially outperforms all other LLMs (basic GPT-3.5-turbo, Bloom and CodeT5). These results highlight the significant advancement of our approach for the text-to-testcase generation task when compared to other baseline LLMs.
The basic usage of ChatGPT may not produce a desirable performance for the text-to-testcase generation task. For CodeT5 and Bloom with the other three scenarios (i.e., No Fine-Tune +Basic P, No Fine-Tune + Imp P, Fine-Tune + Basic P), these models achieve a very low-performing performance (e.g., less than 1% syntax correctness of the generated test cases), meaning that they are not capable to be used in practices. When comparing the BloomFineTune+Imprv.P model and the GPT3.5 NoFineTune+Basic.P model, we found that the basic usage of ChatGPT (GPT-3.5-turbo with no fine-tuning and a basic prompt) may not achieve a desirable performance as it performs worse than the fined-tuned Bloom with the improved prompt design.
Nevertheless, our approach (the GPT3.5 FineTune+Imprv.P model) still performs the best.
RQ2: What is the effect of the fine-tuning step for the LLM-based text-to-testcase generation task?
Results. When fine-tuned the LLMs (GPT-3.5-turbo) with our curated fine-tuned dataset, the performance is improved for all evaluation aspects when compared to the basic usage of ChatGPT. When comparing the models between GPT3.5 NoFine-Tune + Basic.P and GPT3.5 Fine-Tune + Basic.P, we observe an improvement of the syntax correctness by 223%, the requirements alignment by 164%, and the code coverage by 153% (see Figure 5). In terms of syntax correctness, we found that the performance is improved from 12.90% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 41.60% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P.
In terms of requirement alignment, we found that performance improves from 11. 10% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 29.25% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. In terms of code coverage, we found that performance is improved from 10.44% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 26.23% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. These results confirm that the fine-tuning step with our own curated dataset substantially improves the performance of GPT-3.5, suggesting that model fine-tuning is a crucial step when using a large language model like GPT-3.5, which should be considered in the future.
RQ3: What is the effect of the prompting step for the LLM-based text-to-testcase generation task?
Results. With our improved prompts, the performance is improved for all evaluation aspects when compared to the basic usage of ChatGPT. When comparing the models between GPT3.5 NoFine-Tune + Basic.P and GPT3.5 Fine-Tune + Basic.P, we observe an improvement of the syntax correctness by 223%, the requirements alignment by 164%, and the code coverage by 153% (see Figure 5). In terms of syntax correctness, we found that the performance is improved from 12.90% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 28.90% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. In terms of requirement alignment, we found that the performance is improved from 11.10% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 20.25% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P. In terms of code coverage, we found that the performance is improved from 10.44% for the GPT-3.5 model without fine-tuning GPT3.5 NoFine-Tune + Basic.P to 16.43% for the GPT-3.5 model with fine-tuning GPT3.5 Fine-Tune + Basic.P.
These results confirm that the improved prompt substantially improves the performance of GPT-3.5, suggesting that prompting design is a crucial step when using a large language model like GPT-3.5, which should be considered in the future.
RQ4: What types of errors occur in the incorrectly generated test cases?
Results. The majority of the generated test cases by our approach are syntactically correct for 78.5%, while being incorrect for 21.5%. The most common types of errors are assertion errors (11.3%), value errors (2.4%), syntax errors (6.9%), and others (0.9%). Understanding the incorrectly generated test cases by our approach is crucial as it helps researchers gain a better understanding of the limitations of our approach. To do so, we utilized Apache Maven (maven, 2023) as our build tool, and the Surefire Plugin (surefire, 2023) to execute the test cases and to analyze the types of errors during the build process. Figure 6 provides a statistical summary of the error types of the generated test cases. An example of different types of incorrectly generated test cases is also presented in Listing 1. Below, we discuss the different types of errors.
Figure 6: (RQ4)The common types of errors that occur in the generated test cases.
// Assertion error -junit.framework.AssertionFailedError: Escape character should match the default escape character
@Test
public void testGetEscape() {
    CSVFormat format = CSVFormat.DEFAULT;
    char expectedEscape = ’\\’;
    assertEquals("Escape character should match the default escape character", expectedEscape, format.getEscape());}
// Syntax error - java: <identifier> expected, missing ’{’ in the method
@Test
public void testIsSurroundingSpacesIgnored()
    CSVFormat format = CSVFormat.DEFAULT;
    boolean spacesIgnored = format.isSurroundingSpacesIgnored();
    assertTrue("Surrounding spaces should be ignored by default", spacesIgnored);}
// Value error -junit.framework.ComparisonFailure: Line separator should match the default line separator
@Test
public void testGetLineSeparator() {
    CSVFormat format = CSVFormat.DEFAULT;
[…]
    assertEquals("Line separator should match the default line separator", expectedLineSeparator, format.getLineSeparator());}
Listing 1: Types of Error
  • 1.
    Assertion errors (11.3%) occur when an assertion statement (a statement that checks if a condition is true) fails because the condition was false. The results suggest that a noticeable fraction of the test cases had issues with the logical conditions expected to be true during the tests. The assertion error ”Escape character should match the default escape character” indicates a failure in a JUnit test case.
  • 2.
    Syntax errors (6.9%) refer to errors that do not conform to the language rules. As a result, the generated test cases cannot be parsed or compiled. For example, the syntax error in the provided code is due to the missing opening curly brace ’{’ to begin the method body after the method declaration. The code should have an opening brace ’{’ after the method declaration to define the scope of the method and include the test logic inside it.
  • 3.
[…]
  • 4.
    The other (0.9%) category probably includes all other types of errors that do not fit into the aforementioned categories.
6 Threats to Validity
Below, we discuss the most pertinent threats to the validity of our study.
Internal Validity. The main internal validity concerns in our study are two-fold: (i) hyper-parameter optimization of various algorithms and (ii) the potential training data-leakage issue of LLMs.
In our fine-tuning step, we employed a variety of hyperparameter configurations, including adjustments in the number of epochs, batch size, and learning rate. The hyperparameter optimization can significantly improve evaluation outcomes (Bischl et al., 2023).
However, identifying the optimal settings is resource-intensive due to the expansive search space of hyperparameters. We note that our primary objective in this study is not to pinpoint the most effective hyperparameters but to investigate the potential of LLMs in TDD. Thus, while the results achieved with LLMs in our study are promising, they only represent a starting point and have the potential for further improvement through extensive hyperparameter tuning. Nonetheless, we have included detailed information about the hyperparameter settings used in our study within our replication package. The second major concern related to LLMs is the training data leakage (Xia et al., 2023). There is a possibility that GPT-3.5-turbo or other LLMs may have been exposed to the same open-source projects that we used in our study. However, the pre-trained models did not yield satisfactory results, compared to the fine-tuned model and custom prompt results. This shows that the impact of training data leakage, if any, was minimal in our study settings.
Construct Validity. Our study’s main construct validity concern is related to the prompt tuning and robustness (Lester et al., 2021). Given the large scope of subjectivity in prompting and the lack of empirical guidelines on prompt tuning and robustness in SE, one can always further tune and build a more robust prompt to achieve better evaluation results. However, we experimented with several prompt combinations in our journey, from the basic to the custom prompt, in an attempt to mitigate the threats related to prompt tuning.
External Validity The external validity concerns the generalization of our findings. We experimented with five open source Java projects from Github, selected with diversity in mind to improve our external validity. The five projects represent different domains of inputs (string, int, etc.) and sizes and complexity of classes. Further experiments with other open-source projects would help with the generalizability of the results.
7 Conclusion
In this paper, we introduce a Text-to-Testcase generation approach based on a fine-tuned large language model with an effective prompting design. Our results show that the prompting-based fine-tuning GPT3.5 outperforms other baselines. Furthermore, our ablation study demonstrates the substantial performance improvement of the fine-tuning and prompting components of the GPT3.5 model, highlighting the importance of the fine-tuning and prompting components for the text-to-test case generation task. Based on our results, we draw the following recommendations: (1) the basic usage of ChatGPT may not produce a desirable performance for the text-to-testcase generation task; and (2) both fine-tuning and effective prompting design should always be considered when using a large language model.